OpenStudy (anonymous):

Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $123.00 for 3 days and 300 miles, while Mary was charged $216.00 for 5 days and 600 miles. What does Best Rental charge per day, for mileage? Can anyone help me solve this one...I need to build the equations and show the work

OpenStudy (dumbcow):

Let D be daily fee, M be mileage fee Barney: 3D + 300M = 123 mary: 5D + 600M = 216 Its a system of 2 equations solve using substitution or elimination

OpenStudy (dumbcow):

can you solve it from here?

OpenStudy (anonymous):

8D+900M=339 is that right?

OpenStudy (anonymous):

I got: d= 30 m=.11 or 11cents per mile.. good job dumbcow

OpenStudy (dumbcow):

yes, however that doesn't eliminate any variable so it doesn't help

OpenStudy (anonymous):

ok how do I solve?

OpenStudy (dumbcow):

yes thats what i get as well. LCM of 3 and 5 is 15 Multiply top equation by -5 Multiply bottom equation by 3 -15D -1500M = -615 15D +1800M = 648 -------------------- 0 + 300M = 33 M = 33/300 = 0.11

OpenStudy (dumbcow):

Then substitute that value for M back into 1 of the equations to solve for D. 3D +300(0.11) = 123 3D +33=123 3D = 123-33 = 90 D = 90/3 = 30